a tire 0.500m in radius rotates at a constant rate od 200 rev/mon. Find the speed and acceleration of a small stone lodged in the tread of the tire (on its outer edge)
1
Expert's answer
2021-01-31T19:45:57-0500

Given:

"R=0.500\:\rm m;"

"n=200\:\rm rev/min."

The angular velocity


"\omega=2\pi n=2\pi\times 200=1256\:\rm rad/min=20.9\:\rm rad/s."

The linear speed


"v=\omega R=20.9\times 0.500=10.45\:\rm m/s."

The acceleration


"a=\omega^2R=20.9^2\times 0.500=218\:\rm m/s^2."